The profit P, in dollars, gained by selling x computers is modeled by the equation P= -5x^2 + 1000x + 5000. Find the number of computers that must be sold to obtain a profit of $55,000.00?
Using the quadratic equation, I came up with the computer cost being $41.42 then divided $55,000 by that and came out with 1327.86 computers needed to be sold to obtain that profit of $55,000.
Please help.
Diana

Found 2 solutions by stanbon, dfrazzetto:
Answer by stanbon(75887)

The profit P, in dollars, gained by selling x computers is modeled by the equation P= -5x^2 + 1000x + 5000. Find the number of computers that must be sold to obtain a profit of $55,000.00?
Using the quadratic equation, I came up with the computer cost being $41.42 then divided $55,000 by that and came out with 1327.86 computers needed to be sold to obtain that profit of $55,000.
-------------------
-5x^2 + 1000x + 5000 = 55000
Divide both sides by -5 to get
----
x^2 - 200x -1000 = -11000
-----------
x^2 - 200x + 10000 = 0
-----
x = 100
============
Cheers,
Stan H.

Answer by dfrazzetto(283)

P= -5x^2 + 1000x + 5000
55000 = -5x^2 + 1000x + 5000
To use quadratic, we need the left side to be zero:
0 = -5x^2 + 1000x + 5000 - 55000

0 = -5x^2 + 1000x - 50000
Divide by -5:
0 = x^2 - 200x + 10000
factors into 0= (x-100)^2
x=100 computers
CHECK:
P= -5x^2 + 1000x + 5000.
P= -5(100)^2 + 1000(100) + 5000.
P = -50,000 + 100,000 + 50000
P = 55,000√